OpenStudy (anonymous):

Given the equation 5 + x - 12 = x - 7. Part A. Solve the equation 5 + x - 12 = x - 7. In your final answer, be sure to state the solution and include all of your work. Part B. Use the values x = -4, 0, 5 to prove your solution to the equation 5 + x - 12 = x - 7. In your final answer, include all of your calculations.

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

@hartnn lol

OpenStudy (anonymous):

@kryton1212 u gonna help or no

OpenStudy (anonymous):

i got x nothing :O

hartnn (hartnn):

x gets cancelled and we get something like -7 =-7 or 0=0

hartnn (hartnn):

and that is true for ALL VALUES OF X

hartnn (hartnn):

hence the answer is INFINITE solutions

hartnn (hartnn):

thats part A for part B, plug in x =-4 in 5 + x - 12 = x - 7

hartnn (hartnn):

5+4-12 = 4-7 -3 =-3 which is true hence x = -4 is a solution

hartnn (hartnn):

do the same for other values of x

OpenStudy (anonymous):

@hartnn @kryton1212 thanks i need help with more ...

OpenStudy (anonymous):

feel free to ask

OpenStudy (anonymous):

i help nothing lol no need to thanks me haha

OpenStudy (anonymous):

Given the equation 5 + x - 14 = x - 7. Part A. Solve the equation 5 + x - 14 = x - 7. In your final answer, be sure to state the solution and include all of your work. Part B. Use the values x = -2, 0, 3 to prove your solution to the equation 5 + x - 14 = x - 7 . In your final answer, include all of your calculations.

OpenStudy (anonymous):

the approach is the same as what @hartnn ha mentioned few minutes ago

OpenStudy (anonymous):

iknow but this is a different question